OpenStudy (anonymous):

Yuki drove 450 miles in 7 hours. At that rate, how many miles could she drive in 12 hours? Round your answer to the nearest mile

OpenStudy (lgbasallote):

yuki drove 450 miles in 7 hours let's get her speed speed = distance/time speed = 450/7 speed = 6.43 mph to get how far she can go in 12 hours we use distance = speed x time distance = 6.43 x 12

OpenStudy (lgbasallote):

i dont have a calculator so i leave the calculation to you do you get it though?

OpenStudy (shane_b):

\[\frac{450}{7}=\frac{X}{12}\]Solve for x.